BBS.LIST;1 BIBLE.DIR;1 BITNET.DIR;1 BOOKS.DIR;1 CHRIS.DVI;7 CHRIS.LIS;10 CHRIS.SAVE;1 CHRIS.TEX;13 CLASSES.DIR;1 DEVTYPE.INF;1 EX1.TEX;1 EXPENSE.DIR;1 FILENAMES.TEX;1 GIFTS.OUT;1 INVEST.DIR;1 JOB.MEM;1 JOKES.TXT;1 JUDY.DVI;1 JUDY.FOR;1 JUDY.SAVE;1 JUDY.TEX;1 LETTER.RS;1 LOGIN.1986;1 LOGIN.1987;1 LOGIN.TIME;1 LORI.TEX;3 MANAGE.DIR;1 MITES.DIR;1 MWR.COM;1 OLDUN.LIS;1 PHARM304.TEX;13 PHONES.LIS;1 RESMACS.TEX;7 RESUME.DVI;2 RESUME.DVI;1 RESUME.LIS;2 RESUME.LIS;1 RESUME.TEX;8 RESUME.TEX;7 RESUME.TEX;6 SOP.DVI;1 SOP.TEX;1 STANFORD.PURPOSE;1 SWITCH.CLS;1 TEMPLAT.DATA;2 TERMS.LOC;1 TEXT.DIR;1 VAXMACRO.HLP;1 WHO.LIS;1